RC Meeting, Monday, January 10, 2022

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 AM, remote

  • Zoom remote connection information:
 Join ZoomGov Meeting
 https://jlab-org.zoomgov.com/j/1607100885?pwd=ZFBsRTFsYm5Zb0laZ0tyZXNjcEE0QT09
 Meeting ID: 160 710 0885
 Passcode: 595071
 One tap mobile
 +16692545252,,1607100885# US (San Jose)
 +16468287666,,1607100885# US (New York)
 Dial by your location
       +1 669 254 5252 US (San Jose)
       +1 646 828 7666 US (New York)
       +1 669 216 1590 US (San Jose)
       +1 551 285 1373 US
       833 568 8864 US Toll-free
 Meeting ID: 160 710 0885
 Find your local number: https://jlab-org.zoomgov.com/u/ab2CekLXgQ

Useful Links


Announcements

  • Now using Zoom for RC meetings
  • Accelerator restoration in progress

Current Status

Current RC: Dave Gaskell

  • Hall C in Controlled Access
  • Controlled Access today to address some issues with spectrometer cryo:
    • External EPICS signals not updating in PLC
    • SHMS vacuum readback gauge
    • SHMS bender warm return
    • SHMS Q1 leadflow controller
  • DC power also needs to look at small Moller quad power supply (used for beamline optics - won't cycle)
  • Spectrometer magnets off
  • SHMS rotated to 5.5 degrees for start of run (38 degree rotation has been verified)
  • Detector HV on, cosmic runs in progress
  • Beam energy measured over the weekend - good agreement with December measurement: https://logbooks.jlab.org/entry/3968978

Future plans

  • Dave Meekins will condense LH2/LD2 on Tuesday (day) - target operaters will be required from that point
  • Start beam delivery after cooldown

Current Run Plan

  • Start run with carbon hole check to verify beam/target position - may request adjustment to target ladder after this
    • Maybe start with the HMS at larger angle for carbon hole, to avoid possible rotation issues?
  • Begin Q2=3.85, W=3.07, low epsilon data taking
    • SHMS settings: theta = 5.50 degrees, P=6.538 GeV
    • HMS settings: theta = 35.12 degree, P=-1.332 GeV
    • Targets = LH2, dummy, LD2

Notices in Effect for Shift Crew

  • The Raster Coil pair Ax and Ay are off. Raster needs to be at 4x4. Raster gui has been updated - can now request 2x2.
  • With LHe alarms it is good to check the cameras for plumes. If no plume and it is below the LO alarm and above LOLO then monitor with a plot and halog the problem.
  • With LNitrogen alarms are not that critical. Just need to start monitoring the problem with plot and Halog.

Current Issues

  • Vertical beam position on target has changed a little - will ask Dave Meekins to adjust target ladder positions when beam returns
  • There is a rotation issue that has come up if moving HMS between 19 to 25 degrees. It usually gets stuck around this range, please be wary of this issue. We will likely skip settings in this range if this drive controller error cannot be solved. For singles it will not be a big deal to skip settings for HMS at those angles. We need a longer access to address this issue.
  • The Raster Coil pair Ax and Ay are off
  • Diffuser limit at 46uA (Note: is this still valid?)


Crew Chiefs, We have permission to raise the Hall C diffuser raster limit to allow running >46uA to Hall C. This limit is normally imposed when one of the two rasters is not operative. Since the hall was able to double the running raster current to get the same size raster on the target as having both rasters, Keith is allowing us to raise the limit. To get to the setting, open the Hall C diffuser interlocks screen. Go to the bottom right corner and open the Expert screen. Just below and to the right of the FSD bits, raise the "Raster 1 OK Limit" to slightly above the highest current the hall wants to run at. Do not change the Raster 2 limit. Make a log entry with screen shots. Make an ABIL entry, System MPS and an expected removal date of 3 Jan. 2022. Paul

Planned Accesses to the Hall

  • Controlled access today

Planned Accelerator Configuration Changes

  • None

Special Requests

  • None